OpenStudy (anonymous):

establish the identity of (1-sec x / tan x) + (tan x / 1- sec x ) =-2 csc x

OpenStudy (ybarrap):

\[(1-\sec x) / \tan x + \tan x / (1- \sec x ) \] \[u = \tan x/(1-\sec x)\] Equation has form \[1/u + u\] \[=(u^2 + 1)/u\] \[\frac{ (\tan x/ (1-\sec x)) ^2+1 }{\tan x/ (1-\sec x) }\] \[=\frac{ \sin^2 x/(cosx -1)^2 + 1 }{ \tan x/ (1-\sec x) }\] \[=\frac{ \sin^2 x + \cos^2 x - 2\cos x +1 }{ \tan x/ (1-\sec x) }\] \[=\frac{ 2-2cosx }{ sinx/(\cos x(1-1/\cos x) }\] \[=\frac{ \csc x(2 -2\cos x) }{ \cos x -1 }\] \[=\frac{ 2\csc x (1-\cos x) }{ -(1-\cos x) }\] \[=-2\csc x\]

OpenStudy (ybarrap):

i left a few steps out for you to fill in

OpenStudy (anonymous):

yeah solving with the u is a lot easier thanks so much

OpenStudy (ybarrap):

np